Mark is shopping for a new mobile phone. The


probability that someone buys Oppo is 25%,


Huawei is 20%, Vivo is 10%, both Oppo and


Vivo is 5%, both Oppo and Huawei is 7%,


both Vivo and Huawei is 3%, and of the three


brands is 2%. What is the probability that


Mark buys Oppo or Huawei?


1
Expert's answer
2022-02-24T17:49:28-0500


Let "O" denote the event "buy Oppo", "H" denote the event "buy Huawei,"

"V" denote the event "buy Vivo".


"P(O)=0.25, P(H)=0.20, P(V)=0.10"

"P(O\cap V)=0.05, P(O\cap H)=0.07,"

"P(H\cap V)=0.03,P(O\cap H\cap V)=0.02"

The probability that Mark buys Oppo or Huawei is


"P(O\cup H)=P(O)+P(H)-P(O\cap H)"

"=0.25+0.20-0.07=0.38"